A chemist has a bottle of 1% acid solution and a bottle of 5% acid olution. She wants to mix the two solutions to get 100ml of a 4% acid sloution. Write the system of linear equations to determine how many ml of each sloution is required to get 100ml of 4% slution. (Solve the equation and answer the question for five bonus points)
Answer by checkley75(3666)

.04*100=.05x+.01(100-x)
4=.05x+1-.01x
.05x-.01x=4-1
.04x=3
x=3/.04
x=75 ml of 5% solution
100-75=25 ml of 1% solution
proof
.04*100=.05*75+.01*25
4=3.75+.25
4=4